# Stochastic Heating of a Bose-Einstein Condensate

### v1

@@ Abstract
Figures 2(b) and 4(e) further show that the full momentum statistics, namely, the Gaussian distribution of Eq. (2) agrees well with the experimental data in the entire course of time evolution. According to Eq. (4), as the time increases, the variance crosses over from a linear growth: $\mathrm{Var}\,[p_{c,y}]=\Gamma t$ to saturation: $\mathrm{Var}\,[p_{c,y}]=\Gamma/(2\gamma)$ at the time scale $t_{\rm Th}=1/(2\gamma)$. The linear growth suggests that for $t\ll t_{\rm Th}$ the condensate COM undergoes a diffusive motion in the momentum $p_{c,y}$ space. This has a far-reaching consequence: Similar to stochastic heating in plasma physics Fermi (1949); Stix (1964); Hall and Sturrock (1967), because a BEC atom has mean kinetic energy $\mathrm{Var}\,[p_{c,y}]/(2m^{*})$, it gets heated up with a constant rate $\Gamma/(2m^{*})$ by random momentum scattering. For longer times, the cooling comes into play. According to Eq. (1) the cooling rate is $\gamma\mathrm{Var}\,[p_{c,y}]/m^{*}$, which balances the heating rate only if $\mathrm{Var}\,[p_{c,y}]$ attains the saturation value $\Gamma/(2\gamma)$: This implies that the COM reaches thermal equilibrium at $t\sim t_{\rm Th}$, and thus $t_{\rm Th}$ gives the thermalization time. Due to Eq. (2) the entire BEC, a ‘classical’ object, displays a Maxwell-Boltzmann distribution at the kinematic temperature
$T=N_{\rm BEC}\frac{\mathrm{Var}[p_{c,y}(t\gtrsim t_{\rm Th})]}{2k_{B}m^{*}}=\frac{N_{\rm BEC}\Gamma}{4k_{B}\gamma m^{*}},$ (5)
@@
with $k_{B}$ being the Boltzmann constant. Accordingly, for fixed $N_{\rm BEC}$ and $T$, the enhancement of $m^{*}$ leads to the enhancement of the saturation value $\Gamma/(2\gamma)$, consistent with the measured temporal profiles of $\mathrm{Var}[p_{c,y}]$ in Fig. 3(b). As detailed in Ref. sup , we can determine ${\Gamma/(2\gamma)}$ from experimental data and of $m^{*}$ from calculations. Substituting them and the estimation $N_{\rm BEC}\simeq 10^{3}$ into Eq. (5), we obtain $T\simeq 25\,\rm nK$ for all different cases. This is consistent with an alternative estimation based on the trap depth in the gravity direction Wang et al. (2021) and the width of the second band, giving $T\simeq 15\,\rm nK$.
The observed temporal profile of $\mathrm{Var}\,[p_{c,y}]$ resembles dynamical localization Casati et al. (1979); Moore et al. (1995); Chabé et al. (2008) and its many-body variant Toh et al. (2022); Cao et al. (2022); Guo et al. (2023) in driven cold atomic gases. However, there are conceptual differences. Notably, dynamical localization originates from wave interference and mimics Anderson localization in the momentum space Fishman et al. (1982); Tian and Altland (2010), leading to an exponential distribution at equilibrium. In stark contrast, interference plays no role here, and the equilibrium distribution is Maxwell-Boltzmann.
In summary, we have demonstrated that in a BEC interacting with a surrounding thermal cloud and subjected to evaporative cooling induced by atom loss, a sudden quench can lead to an unconventional thermalization process, distinct from the various thermalization mechanisms intensively discussed in previous work D’Alessio et al. (2016); Borgonovi et al. (2016); Nandkishore and Huse (2015); Ueda (2020). Further progress requires a thorough understanding of this interaction on a microscopic level. The present investigations could be extended to systems quenched across the quantum critical point, in order to explore the interplay between the thermalization process described in the present work and an expected Kibble-Zurek mechanism Kibble (1976); Zurek (1985). Finally, in the present work, the stochastic heating arises from Gaussian white noise. In view of the discovery of quantum turbulence in cold atomic gases Navon et al. (2016), it is conceivable that the origin of stochasticity can be more complex leading to richer physics in other non-equilibrium cold atomic systems. This raises the question whether heating by turbulence, which lies at the heart of modern plasma physics Petrosian (2012), has analogies in those systems. Thorough studies of these issues may open a new interdisciplinary research area.
@@
We thank Hui Zhai, Congjun Wu, Shuai Yin, Qi Zhou, Yu Chen, and Masahito Ueda for useful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1404103), NSFC (Grants No. 12274196 and No. 12304289), and funds from Guangdong province (Grants No. 2019QN01X087 and No. 2019ZT08X324); C.T. acknowledges support from NSFC (Grants No. 11925507, No. 12475043 and No. 12047503); S.Z. acknowledges support from HK GRF (Grant No. 17306024), CRF (Grants No. C6009-20G and No. C7012-21G), and a RGC Fellowship Award No. HKU RFS2223-7S03; A.H. acknowledges support from the Cluster of Excellence CUI: Advanced Imaging of Matter of the Deutsche Forschungsgemeinschaft (DFG) - EXC 2056 - project ID 390715994.
@@
Here, the vectors $\mathbf{b}_{j}$ are defined as $\mathbf{b}_{1}=\mathbf{k}_{1}-\mathbf{k}_{2}$, $\mathbf{b}_{2}=\mathbf{k}_{2}-\mathbf{k}_{3}$, and $\mathbf{b}_{3}=\mathbf{k}_{3}-\mathbf{k}_{1}$ with the wave vectors ${\bf k}_{1}=k_{L}(-\sqrt{3}/2,1/2)$, ${\bf k}_{2}=k_{L}(\sqrt{3}/2,1/2)$, and ${\bf k}_{3}=k_{L}(0,-1)$, where $k_{L}=2\pi/\lambda$ and $j=1,2,3$. $\Delta\eta_{j}=2\pi\Delta\nu\Delta L_{j}/c$ determine the difference of the centers of the two triangular lattices. Experimentally, we choose $(\Delta L_{1},\Delta L_{2},\Delta L_{3})\simeq(-6.12,3.06,3.06)\,\rm cm$ by changing the optical paths of the laser beams forming the triangular lattices. Since $\Delta L_{2}=\Delta L_{3}$, tuning $\Delta\nu$ thus changes the separation of the centers of the two triangular lattices along the $x$-axis. In the case of $\Delta\nu=3.263\,$GHz, $V(\mathbf{r})$ becomes a regular hexagonal lattice. For other cases, a deformed hexagonal lattice is realized.
@@
In our experiment, we consider a situation with different final lattice depths among the A and B sites with $(V_{A},V_{B})=(7.32,7.93)\,E_{\rm{R}}$. When the atoms are transferred into the second Bloch band, they mainly populate the shallow $s$-orbitals of the A sites. This leads to an extremely long lifetime. Via tuning $\Delta\nu$, we can significantly change the band dispersion along the $y$ direction, which is shown in Fig. 1(b). In this case, the critical point is estimated to be located around $\Delta\nu_{C}=3.235\,\rm GHz$. To characterize the band dispersion along the $y$ direction, we fit it with $\mathcal{E}=\mathcal{E}_{0}+\frac{1}{2}a\,p_{y}^{2}+\frac{1}{4}b\,p_{y}^{4}$ along the line connecting the $K_{1}$ and $K_{2}$ points. When $\Delta\nu<\Delta\nu_{C}$, $a>0$, while in the opposite case, the band structure shows a double-well dispersion with $a<0$. Focusing on the case with $\Delta\nu<\Delta\nu_{C}$, we numerically calculate the corresponding parameter of $a$ to evaluate the effective mass along the $y$ direction denoted as $m^{*}=1/a$. Details are shown in Fig. S1(a). We approximately choose the overall trapping potential via including the contribution from the dipole trap and the optical lattice along the $y$-direction to be $2\pi\times 50\,\rm Hz$. This enables us to roughly estimate the oscillation period $2\pi/\omega_{0}$, where $\omega_{0}\equiv\sqrt{\kappa/m^{*}}$. Details are shown in Fig. S1(b). Furthermore, using the value of $\gamma$ from experimental data fitting, we caculate $\omega_{1}$ shown in Fig. S1(c). The obtained oscillation period $2\pi/\omega_{1}$ is consistent with experimentally observed oscillation period for the mean value of momentum along the $y$ direction shown in Fig. 3(c).
@@
Figure S2: The estimated temperature for the thermal bath is plotted versus the frequency difference $\Delta\nu$.
@@
$N(t+\delta t)p_{c,y}(t+\delta t)+N_{L}(t)p_{L}(t)$ (S2)
$=$ $N(t)\left(p_{c,y}(t)-\kappa y_{c}\delta t\right),$
@@
$f(p_{c,y};t)$ $=$ $\int{d\eta\over 2\pi}{\rm e}^{i\eta[p_{c,y}-A_{p}{\rm e}^{-\frac{\gamma t}{2}}\sin(\omega_{1}t+\phi_{p})]}$ (S12)
×$\times{\rm e}^{-{\Gamma\over 4\gamma}\{1-{\rm e}^{-\gamma t}[c_{p}^{2}-{c_{p}\over s_{p}}\sin(2\omega_{1}t+\varphi)]\}\eta^{2}},\,\,\,\,\,\,$
where $A_{p}=\sqrt{({1\over 2}\,p_{0}\gamma+\kappa y_{0})^{2}/\omega_{1}^{2}+p_{0}^{2}}$, $\tan(\phi_{p})=-p_{0}\omega_{1}/({1\over 2}\,p_{0}\gamma+\kappa y_{0})$, $\omega_{1}^{2}=\omega_{0}^{2}-\gamma^{2}/4$, $c_{p}=\omega_{0}/\omega_{1}$, and $\tan(\varphi)=1/s_{p}$. Integrating out $\eta$, we obtain Eqs. (2)-(4) in the main text.

### v2

@@ Abstract
Figures 2(b) and 4(e) further show that the full momentum statistics, namely, the Gaussian distribution given by Eq. (2) agrees well with the experimental data throughout the evolution. By Eq. (4), at the time scale $t_{\rm Th}=1/(2\gamma)$ a crossover from $\mathrm{Var}\,[p_{c,y}]=\Gamma t$ to saturation: $\mathrm{Var}\,[p_{c,y}]=\Gamma/(2\gamma)$ occurs. The linear growth suggests that for $t\ll t_{\rm Th}$ the condensate COM undergoes a diffusive motion in the momentum space. Since a BEC atom has mean kinetic energy $\mathrm{Var}\,[p_{c,y}]/(2m^{*})$, it gets heated up with a constant rate $\Gamma/(2m^{*})$ by random momentum scattering. For longer times, the cooling comes into play. According to Eq. (1) the cooling rate is $\gamma\mathrm{Var}\,[p_{c,y}]/m^{*}$, which balances the heating rate only if $\mathrm{Var}\,[p_{c,y}]$ attains the saturation value $\Gamma/(2\gamma)$: This implies that the COM of the BEC reaches thermal equilibrium at $t\sim t_{\rm Th}$ with Eq. (2) being a Maxwell-Boltzmann distribution, as if the BEC as a whole was a classical object. The kinematic temperature is
$T=N_{\rm BEC}\frac{\mathrm{Var}[p_{c,y}(t\gtrsim t_{\rm Th})]}{k_{B}m^{*}}=\frac{N_{\rm BEC}\Gamma}{2k_{B}\gamma m^{*}},$ (6)
@@
with $k_{B}$ being the Boltzmann constant. Accordingly, for fixed $N_{\rm BEC}$ (condensed atom number) and $T$, the enhancement of $m^{*}$ leads to the enhancement of the saturation value $\Gamma/(2\gamma)$, consistent with the measured temporal profiles of $\mathrm{Var}[p_{c,y}]$ in Fig. 3(b). As detailed in Ref. sup , we can determine ${\Gamma/(2\gamma)}$ from experimental data and $m^{*}$ from calculations. Substituting them and the estimation $N_{\rm BEC}\simeq 10^{3}$ into Eq. (6), we find $T$ varies in $35$-$60$ nK for $\Delta\nu$ considered in Fig. 3. On the other hand, we expect that $T$ makes no reference to the direction. In the $z$-direction, it is the order of the trap depth along the gravity direction Wang et al. (2021), which is about $40$ nK. We see the two estimations for $T$ are consistent.
Our thermalization scenario differs from other heating processes in quantum gases. Notably, owing to the strong interplay between dissipation and quantum coherence, the BEC is not destroyed upon quench. Opposite to this, in previous experiments on quenching closed (thus non-dissipative) systems Parker et al. (2013); Clark et al. (2016); Nicklas et al. (2015), the BEC is destroyed and structures such as domain walls are formed. Additionally, at ultralow temperatures, it is conceivable that matter wave effects may result in nontrivial heating processes. Indeed, a profile similar to Fig. 2(d) was experimentally observed in driven cold quantum gases Moore et al. (1995), but that is due to Anderson localization Casati et al. (1979); Fishman et al. (1982); Tian et al. (2004); Tian and Altland (2010) and the stationary distribution is exponential.
Summarizing, we have shown that during quench dynamics a BEC interacting with a surrounding thermal cloud and subjected to evaporative cooling induced by atom loss can display classical Brownian motion in momentum space. The ensuing thermalization of the condensate COM degree of freedom differs from various scenarios for thermalization of quantum systems D’Alessio et al. (2016); Borgonovi et al. (2016); Nandkishore and Huse (2015); Ueda (2020). In view of the discovery of quantum turbulence in cold atomic gases Henn et al. (2009); Navon et al. (2016), the origin of stochasticity can be more complex than Gaussian noise assumed in this work, which may potentially lead to rich new non-equilibrium phenomena in cold atomic systems.
@@
We thank Hui Zhai, Congjun Wu, Shuai Yin, Qi Zhou, Yu Chen, and Masahito Ueda for useful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1404103), NSFC (Grants No. 12274196, No. 12304289, and No. 92476101), and funds from Guangdong province (Grants No. 2019QN01X087 and No. 2019ZT08X324); C.T. acknowledges support from NSFC (Grants No. 11925507, No. 12475043 and No. 12047503); S.Z. acknowledges support from HK GRF (Grant No. 17306024), CRF (Grants No. C6009-20G and No. C7012-21G), and a RGC Fellowship Award No. HKU RFS2223-7S03; A.H. acknowledges support from the Cluster of Excellence CUI: Advanced Imaging of Matter of the Deutsche Forschungsgemeinschaft (DFG) - EXC 2056 - project ID 390715994.
@@
Here, the vectors $\mathbf{b}_{j}$ are defined as $\mathbf{b}_{1}=\mathbf{k}_{1}-\mathbf{k}_{2}$, $\mathbf{b}_{2}=\mathbf{k}_{2}-\mathbf{k}_{3}$, and $\mathbf{b}_{3}=\mathbf{k}_{3}-\mathbf{k}_{1}$ with the wave vectors ${\bf k}_{1}=k_{L}(-\sqrt{3}/2,1/2)$, ${\bf k}_{2}=k_{L}(\sqrt{3}/2,1/2)$, and ${\bf k}_{3}=k_{L}(0,-1)$, where $k_{L}=2\pi/\lambda$ and $j=1,2,3$. $\Delta\eta_{j}=2\pi\Delta\nu\Delta L_{j}/c$ determine the difference of the centers of the two triangular lattices. Experimentally, each set of triangular lattice, A or B, is created by three interfering laser beams. The optical path differences are defined as $\Delta L_{1}=L_{1}-L_{2}$, $\Delta L_{2}=L_{2}-L_{3}$ and $\Delta L_{3}=L_{3}-L_{1}$, where $L_{j}(j=1,2,3)$ represents the optical path of the $j$-th laser beam from the splitting point to the center of the lattice. For more detailed information, we refer to Refs. [17,18]. We choose $(\Delta L_{1},\Delta L_{2},\Delta L_{3})\simeq(-6.12,3.06,3.06)\,\rm cm$ by changing the optical paths of the laser beams forming the triangular lattices. Since $\Delta L_{2}=\Delta L_{3}$, tuning $\Delta\nu$ thus changes the separation of the centers of the two triangular lattices along the $x$-axis. In the case of $\Delta\nu=3.263\,$GHz, $V(\mathbf{r})$ becomes a regular hexagonal lattice. For other cases, a deformed hexagonal lattice is realized.
@@
In our experiment, we consider a situation with different final lattice depths among the A and B sites with $(V_{A},V_{B})=(7.32,7.93)\,E_{\rm{R}}$. When the atoms are transferred into the second Bloch band, they mainly populate the shallow $s$-orbitals of the A sites. This leads to an extremely long lifetime. Via tuning $\Delta\nu$, we can significantly change the band dispersion along the $y$ direction, which is shown in Fig. 1(b). In this case, the critical point is estimated to be located around $\Delta\nu_{c}=3.235\,\rm GHz$. To characterize the band dispersion along the $y$ direction, we fit it with $\mathcal{E}=\mathcal{E}_{0}+\frac{1}{2}a\,p_{y}^{2}+\frac{1}{4}b\,p_{y}^{4}$ along the line connecting the $K_{1}$ and $K_{2}$ points. When $\Delta\nu<\Delta\nu_{c}$, $a>0$, while in the opposite case, the band structure shows a double-well dispersion with $a<0$. Focusing on the case with $\Delta\nu<\Delta\nu_{c}$, we numerically calculate the corresponding parameter of $a$ to evaluate the effective mass along the $y$ direction denoted as $m^{*}=1/a$. Details are shown in Fig. S1(a). While the overall effective trapping frequency due to the lattice potential is difficult to measure directly, we numerically calculate its value by using the measured lattice depth and other relevant laser parameters such as beam waists and intensities. As a result, when the contributions from both the dipole trap and the optical lattice are taken into account, we can determine the overall trapping potential along the $y$-direction approximately, yielding an estimated frequency of $2\pi\times 50\,\rm Hz$. This enables us to roughly estimate the oscillation period $2\pi/\omega_{0}$, where $\omega_{0}\equiv\sqrt{\kappa/m^{*}}=2\pi\times 50\sqrt{m_{\rm Rb}/m^{*}}\,\rm Hz$ and $m_{\rm Rb}$ is the atom mass of 87Rb. Details are shown in Fig. S1(b). Furthermore, using the value of $\gamma$ from experimental data fitting, we caculate $\omega_{1}$ shown in Fig. S1(c). The obtained oscillation period $2\pi/\omega_{1}$ is consistent with experimentally observed oscillation period for the mean value of momentum along the $y$ direction shown in Fig. 3(c).
Note that what is essential to our work is the continuous change in the effective mass $m^{*}$, and the presence of a critical point, at which the band dispersion in the $y$-direction changes from quadratic to quartic. In our experiment, it is convenient to implement such modification of the band dispersion by tuning the frequency difference. For other experimental setups, the same band dispersion modification might be realized by tuning other parameters. When the atom loss-induced cooling is further under controls, similar dynamics is expected to occur.
@@
Comparisons of COM temperature by different methods
On one hand, by using Eq. (6) in the main text we can estimate the kinematic temperature $T$ of the condensate COM. For the frequency difference $\Delta\nu$ ranging from $3.21$ GHz to $3.2325$ GHz, we find that $T$ varies from $35$ nK to $60$ nK (Fig. S2). On the other hand, it is reasonable to expect that the kinetic temperature is independent of the direction. When the thermal equilibrium in the $z$-direction is established, the kinetic temperature is the order of the trap depth along the gravity direction, which is about $40$ nK based on our experimental parameters. This is consistent with that given by Eq. (6) in the main text.
Figure S2: The kinematic temperature for distinct frequency differences $\Delta\nu$ is estimated by using Eq. (6) in the main text.
Momentum fluctuations in $\boldsymbol{x}$-direction
In Fig. S3 we show the experimental results for the fluctuations of the $x$-component of the COM momentum of BEC, denoted as $p_{c,x}$. Fig. S3(a) is the temporal profile of the mean and Fig. S3(b) the variance. The experimental protocol and parameters are exactly the same as those in Fig. 2. We see clearly that fluctuations in the $x$-direction are smaller than those in the $y$-direction by over two orders of magnitude.
Figure S3: The measurement of the time evolution of the mean (a) and variance of the BEC COM momentum in the $x$ direction. The experimental protocol and parameters are the same as those in Fig. 2.
Time evolution of atom number in different bands
Band mapping analysis allows us to detect the atom number in the first and the second Brillouin zone (BZ). In Fig. S4 we plot the experimental results for the time evolution of the number of thermal atoms in the first and second BZ, and also atom number of BEC in the second BZ.
Figure S4: The time evolution of the number of thermal atoms in the first and second BZ, as well as the atom number of BEC in the second BZ. $\Delta\nu=3.2275$ GHz.
Determination of error bars
In our data analysis of measured $p_{c,y}$, at each time we fit over $300$ measured data points to a Gaussian distribution of the form: $f(p_{c,y})={1\over\sqrt{2\pi\sigma^{2}}}e^{-(p_{c,y}-\mu)^{2}/(2\sigma^{2})}$ using the fitdist function of MATLAB, which employs maximum likelihood estimation. Then $\mu$ and $\sigma^{2}$ give the mean and the variance of $p_{c,y}$, respectively. The error bars of them represent $68\%$ confidence intervals obtained from Gaussian fitting. They were computed via MATLAB’s paramci function, based on Student’s t-test and Chi-squared test.
Comparison between TOF and band-mapping data
To compare the data obtained from the TOF and band-mapping techniques, in Fig. S5 we have plotted Fig. 4(d) alongside the data (black dots) from Fig. 3(b), both of which correspond to $\Delta\nu=3.2275\,\rm GHz$. The two data sets are in good agreement, which supports the conclusion that the two measurement methods yield consistent results.
Figure S5: The time evolution of the variance of $p_{c,y}$ obtained from TOF and band-mapping measurements at $\Delta\nu=3.2275$ GHz.
@@
$N(t+\delta t)p_{c,y}(t+\delta t)+N_{L}(t)p_{L}(t)=N(t)\left(p_{c,y}(t)-\kappa y_{c}\delta t\right),$ (S2)
@@
$f(p_{c,y};t)=\int{d\eta\over 2\pi}{\rm e}^{i\eta[p_{c,y}-F_{1}(y_{0},p_{0};t)]}{\rm e}^{-F_{2}(t)\eta^{2}},\,\,\,\,\,\,$ (S12)
where
$F_{1}(y_{0},p_{0};t)$ $=$ ${{\rm e}^{-\frac{\gamma t}{2}}\over\omega_{1}}\left[({1\over 2}\,p_{0}\gamma+\kappa y_{0})\sin(\omega_{1}t)-p_{0}\omega_{1}\cos(\omega_{1}t)\right],$ (S13)
$F_{2}(t)$ $=$ $\frac{\Gamma}{4\gamma}\left(1-{\rm e}^{-\gamma t}[c_{p}^{2}-\frac{c_{p}\gamma}{2\omega_{1}}\sin(2\omega_{1}t+\varphi)]\right),$ (S14)
where $\omega_{1}^{2}=\omega_{0}^{2}-\gamma^{2}/4$, $c_{p}=\omega_{0}/\omega_{1}$, and $\tan(\varphi)=1/s_{p}$. Integrating out $\eta$, we obtain
$f(p_{c,y};t)=\frac{1}{\sqrt{4\pi F_{2}(t)}}\exp\left(-\frac{(p_{c,y}-F_{1}(y_{0},p_{0};t))^{2}}{4F_{2}(t)}\right).$ (S15)
Let us introduce two functions defined as
$A_{p}(y,p)\equiv\sqrt{({1\over 2}\,p\gamma+\kappa y)^{2}/\omega_{1}^{2}+p^{2}},\quad\phi_{p}(y,p)=-\arctan\frac{p\omega_{1}}{{1\over 2}\,p\gamma+\kappa y}.$ (S16)
Then we can rewrite $F$ as
$F_{1}(y_{0},p_{0};t)=A_{p}(y_{0},p_{0})\,{\rm e}^{-\frac{\gamma t}{2}}\sin[\omega_{1}t+\phi_{p}(y_{0},p_{0})].$ (S17)
This implies that $A_{p}(y_{0},p_{0})$ and $\phi_{p}(y_{0},p_{0})$ give respectively the amplitude and the initial phase of a damped (monochromatic) oscillation. For the distribution (S15), the first and second moments are
[…]
from which Eqs. (2)-(4) in the main text follow immediately.
Note that $F_{1}$ depends on $p_{0},\,y_{0}$, while $F_{2}$ is independent of $p_{0},\,y_{0}$. Therefore, in each measurement, since the initial phase coordinates $p_{0},\,y_{0}$ are fixed, the time evolution of the Gaussian distribution (S15) is completely fixed. At measurement time $t$, the mean, i.e. the center of distribution $F_{1}$, depends on $p_{0},\,y_{0}$, whereas the variance, i.e. $F_{2}$, does not. Repeating the measurements, we obtain an ensemble of temporal profiles ${\rm E}\,[p_{c,y}]$, which (at given time) fluctuates from measurement to measurement. Let us perform the ensemble average (denoted as $\overline{\cdots}$). Since by Eq. (S13) the dependence of $F_{1}$ on $p_{0},\,y_{0}$ is linear, and all parameters $\gamma,\,\kappa,\,\omega_{1}$ are independent of $p_{0},\,y_{0}$, we have
$\overline{F_{1}(y_{0},p_{0};t)}$ $=$ ${{\rm e}^{-\frac{\gamma t}{2}}\over\omega_{1}}\left[({1\over 2}\,\overline{p_{0}}\gamma+\kappa\overline{y_{0}})\sin(\omega_{1}t)-\overline{p_{0}}\omega_{1}\cos(\omega_{1}t)\right],$ (S20)
$=$ $A_{p}(\overline{y_{0}},\overline{p_{0}})\,{\rm e}^{-\frac{\gamma t}{2}}\sin[\omega_{1}t+\phi_{p}(\overline{y_{0}},\overline{p_{0}})].$
This expression differs from Eq. (S18) only in that in $A_{p}$ and $\phi_{p}$ the arguments $(y_{0},p_{0})$ are replaced by $(\overline{y_{0}},\overline{p_{0}})$. Therefore, we use the same symbol ${\rm E}\,[p_{c,y}]$ for the temperal profiles of ${\rm E}\,[p_{c,y}]$ with and without averaged over reapeated measurements. Eq. (S20) gives the theoretical curve in Fig. 2(c) and Fig. 3(a) of the main text.